1.       Assume that you are working in one of the profit marking firm in town as a research and planning expert. Your organization faces a cost function of C(Q) = 3Q2 + 10Q + 50; and the demand function for the product the firm produces Q = 125 - 5P. If your firm is going to a pricing decision which enable it to maximize its profit. How much price you are expected to advice for the decision maker?

2.      A firm faces a non-linear demand function of P = 250 - 3Q13 - Q12 + Q22  + Q1 - Q2 for two products is the same. Compute the marginal revenue for each product. 


1
Expert's answer
2021-04-08T05:29:20-0400

1.

product demand equation

"Q=125-5P"

"\therefore P=25-0.2Q"

"TC=3Q^2 + 100Q+ 50"

"MC=6Q+100"

"TR=Q(25-0.2Q)"

"=25Q-0.2Q^2"

"MR=25-0.4Q"

To find Q at which MR=MC

"25-0.4Q=6Q+100"

"-75=6.4Q"

"Q=-11.71875"

The  negative output  means that there is slack in the economy due to weak demand.


Price associated with Q=-11.71875

"P=25-0.2(-11.71875)"

"P=22.65625"

I would advice price to be set at 22.65625


2.

"MR=\frac{\Delta TR}{\Delta Q}"

Marginal product for Q1

"\frac{\Delta TR_1}{Q_1}=\frac{250Q_2-3Q_1 ^3Q_2-Q_1^2Q_2+Q_2^3+Q_1Q_2-Q_2^2}{\Delta Q_1}"


"=-9Q_2Q_1^2-2Q_2Q_1+3Q_2^2+Q_1-2Q_2"


Marginal product for Q2


"\frac{\Delta TR_2}{Q_2}=\frac{250Q_2-3Q_1 ^3Q_2-Q_1^2Q_2+Q_2^3+Q_1Q_2-Q_2^2}{\Delta Q_2}"


"=250-3Q_1^3-Q_1^2+3Q_2^2+Q_1-2Q_2"